# UniPart: Part-Level 3D Generation with Unified 3D Geom–Seg Latents

### v1

@@ 4.1 Implementation Details
We construct a curated dataset by integrating multiple public sources [objaverseXL, objaverse, collins2022abodatasetbenchmarksrealworld], yielding 302K objects (184K + 118K) with part-level segmentations. The 118K subset comprises exclusively watertight meshes. Part annotations are derived from mesh connectivity, with each object containing more than 1 part. To ensure topological consistency for SDF computation, all meshes are remeshed using a winding-number-based method [Barill:FW:2018]. Part labels are then propagated to the remeshed geometry by assigning each new face the label of its nearest original face.
@@
We construct a test set of 100 objects with part-level annotations, each including rendered images and ground-truth part geometries. Since there is no reliable correspondence between generated parts and ground-truth parts, we concatenate all parts into a single unified mesh for evaluation to enable a holistic assessment of geometric fidelity. We report Chamfer Distance (CD) and F-Score at thresholds of 0.05 and 0.10 to evaluate fine- and coarse-level geometric alignment, respectively. All meshes are normalized to [1,1][-1,1], and the best score across rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ is selected to ensure a pose-agnostic comparison. We implement five baseline methods for comparison: (i) PartPacker [tang2025partpacker], (ii) PartCrafter [lin2025partcrafter], (iii) OmniPart [yang2025omnipart], (iv) HoloPart [yang2025holopart], and (v) X-Part [yan2025xpart]. Among them, methods (i, ii, iii) accept single-image input for generation, the same as our approach. Methods (iv, v) require mesh input, so we generate the holistic object mesh using Huanyuan3D-2.1 [hunyuan3d2025hunyuan3d21], the same as ours, to enable fair comparison. Note that all baseline methods are implemented with the open-sourced version for evaluation.
@@
For the Geom-Seg VAE and global-level shape generation, we build upon Hunyuan3D-2.1 [hunyuan3d2025hunyuan3d21], using classifier-free guidance (CFG) with a dropout rate of 0.1 and the AdamW optimizer with a weight decay of 0.01. For part-level generation, we initialize the weights from Hunyuan3D-2.1 and fine-tune with the same AdamW settings. All training is conducted on 8 NVIDIA A800 GPUs (80GB) with a batch size of 32. More implementation details are included in the supplementary materials.
@@ 4.2 Experiment Results
We first present the reconstruction performance of our proposed Geom-Seg VAE, as shown in Fig. 2. Both the reconstructed geometry and the inferred part-level latent segmentation demonstrate high fidelity to the input, preserving overall shape and part details while maintaining coherent part assignments across the object. We also provide results from Hunyuan3D-2.1 [hunyuan3d2025hunyuan3d21] as a reference for the reconstructed geometry. These reference reconstructions offer a baseline for comparison, illustrating that our Geom-Seg VAE achieves accurate part-level segmentation while maintaining high-quality geometry. As an intermediate product, the whole-object geometry and part latent segmentation actually provide a solid support for the part-level diffusion at the second stage.
Figure 2: Reconstruction results of our Geom-Seg VAE. (a) Input mesh with part segmentation masks; (b) Reconstructed geometry; (c) Reconstructed part latent segmentation visualization. (d) Reconstructed geometry of Hunyuan3D-2.1 [hunyuan3d2025hunyuan3d21] for reference. Better zoom in for details.
@@
We present both quantitative and qualitative evaluations of the final part-level generation produced by UniPart, and compare against baseline methods, including HoloPart [yang2025holopart], OmniPart [yang2025omnipart], PartCrafter [lin2025partcrafter], PartPacker [tang2025partpacker], and XPart [yan2025xpart]. Qualitative results are provided in Tab. 2. Our UniPart outperforms all other competitors consistently on all metrics (i.e., CD and F-Score with two distinct thresholds), which shows our method produces final objects with higher geometric quality and better alignment to the ground truth. We further provide qualitative results in Fig. 4, which illustrates that our UniPart can generate more reasonable part segmentations and higher-quality part geometries. Some other part-level generations are visualized in Fig. 5. More visual results are shown in the teaser figure and supplementary materials.
@@
HoloPart [yang2025holopart] 0.1492 0.5208 0.7450
OmniPart [yang2025omnipart] 0.1453 0.5273 0.7656
PartCrafter [lin2025partcrafter] 0.1778 0.4749 0.7120
PartPacker [tang2025partpacker] 0.1654 0.4715 0.7226
XPart [yan2025xpart] 0.1533 0.5242 0.7523
@@
Our UniPart actually conducts part segmentation during generation, so the segmentation performance can not be directly evaluated on existing 3D part segmentation benchmarks. Therefore, we specifically construct an evaluation set with 50 object meshes generated by our whole-level DiT (decode the global-geometry latent with Geom. decoder). A group of ten human annotators is invited to manually annotate the part segmentation on these meshes, also given the input image as a reference. The collected annotations are further double-checked and finely corrected by annotators in the second round to guarantee high quality. As a result, we obtain 536 high-quality part annotations from the 50 object meshes. Based on them, we evaluate the part segmentation results of UniPart with existing SOTA baseline methods, including SAMesh [tang2024samesh], PartField [liu2025partfield], and P3-SAM [ma2025p3sam]. For a direct comparison with ours, which only contains 4,096 latent vectors as part segmentation, we uniformly sample 4,096 points from each segmented point cloud or mesh surface of other methods to compute the numerical results. For quantitative results, we adopt mean IoU (mIoU) as the evaluation metric, with results shown in Tab. 2. We found that when the geometry pattern on the mesh surface is not significant enough to support feed-forward-style part segmentation, our diffusion-based generative segmentation can still achieve robust performance, thanks to the part knowledge gained in the generative geometry learning process.
@@
Method SAMesh [tang2024samesh] PartField [liu2025partfield] P3-SAM [ma2025p3sam] UniPart (Ours)
@@ 4.3 Ablation Studies
In our Part-level DiT design, we feed global geometry information as a condition to enable whole-aware part generation. A visual comparison of not using the global geometry condition is shown in Fig. 7. Compared with the full version, removing the global geometry condition influences the seamlessness of the composition. Besides, we also provide a comparison between using latent-form conditions and using classic point-form conditions. In previous methods like X-Part [yan2025xpart] and HoloPart [yang2025holopart], they sample dense points from the segmented object mesh and further train feature extraction layers for geometry encoding, which actually introduces (i) information loss in sampling and encoding; (ii) a domain gap with latent vectors to denoise; and (iii) extra training burden. As shown in Fig. 7, using point-form conditions results in geometry degradation at a fine-grained level, especially for some small-scale parts.

### v2

@@ 4.1 Implementation Details
We construct a curated dataset by integrating multiple public sources [9, 10, 6], yielding 302K objects (184K + 118K) with part-level segmentations. The 118K subset comprises exclusively watertight meshes. Part annotations are derived from mesh connectivity, with each object containing more than 1 part. To ensure topological consistency for SDF computation, all meshes are remeshed using a winding-number-based method [2]. Part labels are then propagated to the remeshed geometry by assigning each new face the label of its nearest original face.
@@
We construct a test set of 100 objects with part-level annotations, each including rendered images and ground-truth part geometries. Since there is no reliable correspondence between generated parts and ground-truth parts, we concatenate all parts into a single unified mesh for evaluation to enable a holistic assessment of geometric fidelity. We report Chamfer Distance (CD) and F-Score at thresholds of 0.05 and 0.10 to evaluate fine- and coarse-level geometric alignment, respectively. All meshes are normalized to [1,1][-1,1], and the best score across rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ is selected to ensure a pose-agnostic comparison. We implement five baseline methods for comparison: (i) PartPacker [31], (ii) PartCrafter [18], (iii) OmniPart [40], (iv) HoloPart [38], and (v) X-Part [37]. Among them, methods (i, ii, iii) accept single-image input for generation, the same as our approach. Methods (iv, v) require mesh input, so we generate the holistic object mesh using Huanyuan3D-2.1 [32], the same as ours, to enable fair comparison. Note that all baseline methods are implemented with the open-sourced version for evaluation.
@@
For the Geom-Seg VAE and global-level shape generation, we build upon Hunyuan3D-2.1 [32], using classifier-free guidance (CFG) with a dropout rate of 0.1 and the AdamW optimizer with a weight decay of 0.01. For part-level generation, we initialize the weights from Hunyuan3D-2.1 and fine-tune with the same AdamW settings. All training is conducted on 8 NVIDIA A800 GPUs (80GB) with a batch size of 32. More implementation details are included in the supplementary materials.
@@ 4.2 Experiment Results
We first present the reconstruction performance of our proposed Geom-Seg VAE, as shown in Fig. 2. Both the reconstructed geometry and the inferred part-level latent segmentation demonstrate high fidelity to the input, preserving overall shape and part details while maintaining coherent part assignments across the object. We also provide results from Hunyuan3D-2.1 [32] as a reference for the reconstructed geometry. These reference reconstructions offer a baseline for comparison, illustrating that our Geom-Seg VAE achieves accurate part-level segmentation while maintaining high-quality geometry. As an intermediate product, the whole-object geometry and part latent segmentation actually provide a solid support for the part-level diffusion at the second stage.
Figure 2: Reconstruction results of our Geom-Seg VAE. (a) Input mesh with part segmentation masks; (b) Reconstructed geometry; (c) Reconstructed part latent segmentation visualization. (d) Reconstructed geometry of Hunyuan3D-2.1 [32] for reference. Better zoom in for details.
@@
We present both quantitative and qualitative evaluations of the final part-level generation produced by UniPart, and compare against baseline methods, including HoloPart [38], OmniPart [40], PartCrafter [18], PartPacker [31], and XPart [37]. Qualitative results are provided in Tab. 2. Our UniPart outperforms all other competitors consistently on all metrics (i.e., CD and F-Score with two distinct thresholds), which shows our method produces final objects with higher geometric quality and better alignment to the ground truth. We further provide qualitative results in Fig. 4, which illustrates that our UniPart can generate more reasonable part segmentations and higher-quality part geometries. Some other part-level generations are visualized in Fig. 5. More visual results are shown in the teaser figure and supplementary materials.
@@
HoloPart [38] 0.1492 0.5208 0.7450
OmniPart [40] 0.1453 0.5273 0.7656
PartCrafter [18] 0.1778 0.4749 0.7120
PartPacker [31] 0.1654 0.4715 0.7226
XPart [37] 0.1533 0.5242 0.7523
@@
Our UniPart actually conducts part segmentation during generation, so the segmentation performance can not be directly evaluated on existing 3D part segmentation benchmarks. Therefore, we specifically construct an evaluation set with 50 object meshes generated by our whole-level DiT (decode the global-geometry latent with Geom. decoder). A group of ten human annotators is invited to manually annotate the part segmentation on these meshes, also given the input image as a reference. The collected annotations are further double-checked and finely corrected by annotators in the second round to guarantee high quality. As a result, we obtain 536 high-quality part annotations from the 50 object meshes. Based on them, we evaluate the part segmentation results of UniPart with existing SOTA baseline methods, including SAMesh [30], PartField [20], and P3-SAM [23]. For a direct comparison with ours, which only contains 4,096 latent vectors as part segmentation, we uniformly sample 4,096 points from each segmented point cloud or mesh surface of other methods to compute the numerical results. For quantitative results, we adopt mean IoU (mIoU) as the evaluation metric, with results shown in Tab. 2. We found that when the geometry pattern on the mesh surface is not significant enough to support feed-forward-style part segmentation, our diffusion-based generative segmentation can still achieve robust performance, thanks to the part knowledge gained in the generative geometry learning process.
@@
Method SAMesh [30] PartField [20] P3-SAM [23] UniPart (Ours)
@@ 4.3 Ablation Studies
In our Part-level DiT design, we feed global geometry information as a condition to enable whole-aware part generation. A visual comparison of not using the global geometry condition is shown in Fig. 7. Compared with the full version, removing the global geometry condition influences the seamlessness of the composition. Besides, we also provide a comparison between using latent-form conditions and using classic point-form conditions. In previous methods like X-Part [37] and HoloPart [38], they sample dense points from the segmented object mesh and further train feature extraction layers for geometry encoding, which actually introduces (i) information loss in sampling and encoding; (ii) a domain gap with latent vectors to denoise; and (iii) extra training burden. As shown in Fig. 7, using point-form conditions results in geometry degradation at a fine-grained level, especially for some small-scale parts.